OpenStudy (anonymous):

The cost of renting a bicycle is $9 for the first hour plus another $4 for each additional hour. Which of the following represents this situation, where C is the total cost of renting a bicycle for h hours?

OpenStudy (anonymous):

I believe you answer would be 9+4x=c 9 is how much the cost for the first hour 4x is the 4 X the numbers of additional hours add C is how much the cost is all you need to the is the x plug in the x to the problem and you should get your answer I will give you an example let say he spent an hour over so that would be 4 more dollars he would have to pay now you right out your problem 9+4x=c we said he only went 1 hour over so we will plug 1 in for x 9+4X1=c now we are going to solve fo 4X1wich is 4 now you problem should look like this 9+4=c add 9+4 now i will looklike this 13=13 now check you answer 13-4=9 is this right yes it is so you woud do 9+4x=c to find what the total cost is hope this helps